OpenStudy (anonymous):

Solve for x. Round your answer to 2 decimal places.

OpenStudy (anonymous):



OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

i know i use sin..

OpenStudy (zzr0ck3r):

\(\frac{\sin(30)}{16}=\frac{\sin(90)}{x}\)

OpenStudy (zzr0ck3r):

solve for x

OpenStudy (anonymous):

32

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

ok, because i was confused if it should be 16sin(30) or 30sin(16) but i guess it was neither :p @zzr0ck3r

OpenStudy (zzr0ck3r):

SOH CAH TOA sin(30) = 16/x thats another way to think about it

OpenStudy (anonymous):

ohhh ok, thank you :)